OpenStudy (trojanpoem):

Find the limit if they exist: f(x) = \[\frac{ asin(x) - xsin(a) }{ x - a } x < a\] f(x) = \[\frac{ \sin(x) - \sin(a) }{ x - a } x > a\]

OpenStudy (yolo_king):

1. Doesn't exist 2. Doesn't exist

OpenStudy (trojanpoem):

Steps ?

OpenStudy (anonymous):

Are these two separate functions, or is it \[f(x)=\begin{cases}\dfrac{a\sin x-x\sin a}{x-a}&\text{for }x<a\\\\ \dfrac{\sin x-\sin a}{x-a}&\text{for }x>a\end{cases}\]

OpenStudy (trojanpoem):

It's like you typed it.

OpenStudy (anonymous):

Also is \(x\to a\) from a particular direction? Or are you asked to find the one-sided limits only?

OpenStudy (yolo_king):

Two separate functions I believe.

OpenStudy (trojanpoem):

Not separated bro.

OpenStudy (yolo_king):

Oh its not :o

OpenStudy (trojanpoem):

You'ill have to find the limit from both sides so as to compare the values if equal it exists

OpenStudy (trojanpoem):

Can you solve it ?

OpenStudy (yolo_king):

Sorry I have no idea, I tried.

OpenStudy (anonymous):

The right-sided limit is fairly simple. Recall the definition of the derivative: \[f'(c)=\lim_{h\to0}\frac{f(c+h)-f(c)}{h}=\lim_{x\to c}\frac{f(x)-f(c)}{x-c}\] This means that \[\large\lim_{x\to a^+}f(x)=\lim_{x\to a^+}\frac{\sin x-\sin a}{x-a}=g'(a)\] where \(g(x)=\sin x\). Hence the right-sided limit is \(\cos a\), since \(\dfrac{d}{dx}\sin x=\cos x\).

OpenStudy (anonymous):

For the left-sided limit, you have \[\begin{align*}\frac{a\sin x-x\sin a}{x-a}&=\frac{a\sin x-(x-a+a)\sin a}{x-a}\\\\ &=\frac{a\sin x-a\sin a-(x-a)\sin a}{x-a}\\\\ &=a\left(\frac{\sin x-\sin a}{x-a}\right)-\sin a \end{align*}\] As before, you'll find that the fraction approaches \(\cos a\), and so the right-sided limit is \(a\cos a-\sin a\).

OpenStudy (trojanpoem):

Cool dude , i really like when you + a and - it again ! But I have a small difficulty to understand from "This mean that to" in the first comment the second one is understandable

OpenStudy (anonymous):

I think what's confusing you is that I'm using \(f\) to mean two different things. The first \(f\), in the definition of the derivative, is a general differentiable function. The second one is the piecewise function given above. Let me write it another way. \[F'(c)=\lim_{h\to0}\frac{F(c+h)-F(c)}{h}=\lim_{x\to c}\frac{F(x)-F(c)}{x-c}\] Denote the piecewise function as \(f\), then \[\lim_{x\to a^+}f(x)=\lim_{x\to a^+}\frac{\sin x-\sin a}{x-a}=\lim_{x\to a^+}\frac{F(x)-F(a)}{x-a}=F'(a)\] Then, since \(F(x)=\sin x\) and \(F'(x)=\cos x\), you get \[\lim_{x\to a^+}f(x)=\lim_{x\to a^+}\frac{\sin x-\sin a}{x-a}=\lim_{x\to a^+}\frac{F(x)-F(a)}{x-a}=\cos a\]

OpenStudy (trojanpoem):

Thanks ! fantastic, really a masterpiece. Could you help me with another one , if you don't mind ?